# Warming-driven rise in soil moisture entropy signals destabilization of the Asian Water Tower

### v1

@@ Abstract
The Tibetan Plateau (TP), known as the “Asian Water Tower,” is currently undergoing a rapid wetting trend. While this moisture increase is often viewed as beneficial for water availability, it remains unclear whether the hydrological system itself is becoming more resilient or drifting toward instability. Here we apply an entropy-based framework to quantify the changing structural organization of the TP’s soil moisture system. We show that from 2000 to 2024, regional wetting has driven a long-term decline in entropy, reflecting an increase in system order and stability due to enhanced hydrological buffering capacity. This stability is modulated by the El Niño–Southern Oscillation (ENSO), which regulates regional heterogeneity via a distinct spatial dipole. Crucially, however, CMIP6 climate projections reveal an alarming reversal: future warming triggers a rise in entropy. This transition signals a loss of systemic resilience, characterized by intensified spatial disorder and potential abrupt regime shifts by the mid-century. Our findings suggest that while current wetting provides a stabilizing buffer, continued warming is projected to amplify spatial heterogeneity, thereby destabilizing the Asian Water Tower, with significant risks for downstream water security.
@@ 1 Introduction
A further challenge lies in the lack of a theoretical framework capable of assessing the systemic behavior of SM beyond localized or pixel-based variability. SM dynamics arise from nonlinear interactions among Pp, ET, snowmelt and vegetation feedbacks, and are additionally influenced by stochastic climate variability [seneviratne2010investigating]. Under external perturbations such as sustained warming, these coupled processes may trigger amplifying mechanisms, such as plant-soil moisture positive feedback [hu2024plant], potentially pushing the system toward alternative regimes when stability decreases. Capturing such large-scale behavior requires a metric that goes beyond mean trends to quantify organization, heterogeneity and resilience. Entropy provides precisely such a diagnostic lens. By quantifying the degree of disorder, entropy reveals whether variability is concentrated in a few dominant modes (order) or dispersed across many (disorder), offering a physically interpretable measure of systemic structure and early-warning signals of abrupt transitions [vranken2015review, tirabassi2023entropy, meng2020complexity]. Here, we leverage recent advances in statistical physics to estimate entropy directly from observational and model data using an eigen microstate approach [hu2019condensation, liu2022renormalization, ma2024increased, zhang2024eigen, wang2024holistic, xie2025ecosystem]. This method extracts dominant system states to characterizes their probability distribution [Liu2025Phase], offering a tractable framework for tracking SM disorder under both observed variability and projected climate warming.
@@ 2.1 Significant wetting over the TP during 2000–2024
Considering the inconsistencies across current SM datasets, we first identify the dataset most suitable for our analysis of the spatiotemporal dynamics of SM across the TP. We develop a robust evaluation framework that jointly considers temporal correlation, spectral characteristics, and autocorrelation structure, using in situ observations at 5 cm depth from the Tibet-Obs network, which comprises Maqu, Naqu, Ali, and Shiquanhe subregions (Supplementary Fig. 1). The evaluation incorporates three complementary metrics designed to assess different aspects of temporal fidelity. (1) temporal correlation ($R_{T}$; see Methods), which quantifies the degree of synchronicity between the evaluated products and site-level measurements. A high $R_{T}$ indicates strong agreement in the timing and fluctuation patterns of SM variability. (2) spectral correlation ($R_{F}$), calculated from the Fourier transforms of evaluated products and observational time series. $R_{F}$ assesses the preservation of frequency-dependent variability, ensuring the dataset reproduces not only first-order trends but also higher-order dynamical structures embedded in the observed record. (3) autocorrelation sequence similarity ($R_{ACF}$), based on lagged autocorrelation coefficients from AC1 to AC10. $R_{ACF}$ captures the persistence and memory structure of SM dynamics, with higher similarity indicating a better capability of the evaluated product to replicate the temporal continuity and internal variability of real-world processes.
@@ 2.2 ENSO-Driven Modulation of SM Entropy on the TP
A broadly increasing trend in SM across the TP is reshaping regional hydrological dynamics by enhancing soil water storage, altering the partitioning of ET, and modifying runoff generation. These changes affect both the timing and magnitude of water availability across the Plateau and its downstream basins, with important implications for the stability of the regional water cycle. To evaluate soil water stability, we use eigen microstate entropy as a diagnostic of system disorder. This metric quantifies how SM variability is distributed across dominant spatial-temporal modes: lower entropy indicates a state that governed by a few coherent modes, whereas higher entropy reflects a more disordered state lacking clear organization (see Methods). Applying this framework to the TP’s SM reveals a distinct long-term reorganization and clear teleconnections with ENSO events.
@@ 2.4 Projected drying and increasing disorder of the SM system over the TP
Rapid warming over the TP is expected to elicit nonlinear responses in the SM system, potentially triggering abrupt regional shifts under future climate scenarios [liu2023teleconnections, terpstra2025assessment, zaitchik2023wetting, qiao2023soil]. Projected increases in ENSO frequency and amplitude [cai2021changing] may further destabilize SM patterns through the teleconnections described above. Moreover, Soil moisture–atmosphere coupling over the TP may propagate via Rossby wave pathways, potentially modulating remote climate extremes, including heatwaves over Europe and East Asia [jiang2024dry, qiao2023soil], highlighting the importance of assessing projected TP soil moisture changes under warming scenarios. In complex systems, rising entropy has been proposed as an early-warning signal of reduced stability, preceding abrupt transitions across diverse contexts, from theoretical models to biological processes [tirabassi2023entropy, vranken2015review]. To illustrate this, we consider a Local Facilitation Cellular Automaton (LFCA) model [kefi2007local], which undergoes a bifurcation-like shift when a control parameter changes. Unlike the one-dimensional Ornstein-Uhlenbeck process often used in climate tipping point studies [smith2023reliability], the LFCA model is inherently high-dimensional, allowing it to capture interactions among multiple components. This property makes it more representative of real-world complexity and well suited to eco-hydrological processes such as vegetation-wetness interactions [tirabassi2023entropy]. As shown in Fig. 5a, system state declines as the control parameter weakens, while the entropy of system state ($S_{\mathrm{LFCA}}$) rises in advance of the threshold, signaling the approach of an abrupt transition.
@@ 3 Discussion and Conclusion
Rising entropy and enhanced spatial heterogeneity across the TP suggest that water availability will become increasingly uneven. This may have cascading implications for major downstream river basins in South and East Asia, which collectively support nearly two billion people. Negative spring SM anomalies in the western TP, for instance, have been linked to extreme summer Pp in the Yangtze River Basin, intensifying downstream flood risk [zhu2023diagnosing]. At the same time, SM-driven circulation anomalies across the Plateau, reinforced by strong land–atmosphere feedbacks, are identified as key drivers of concurrent extremes, including East Asian heatwaves and South Asian floods [he2023extremely]. Increasing TP instability may therefore pose significant threats to agricultural production, hydropower generation, and water security across densely populated regions [yao2022imbalance, cui2023non]. The emergence of regime-shift behavior under high-emission scenarios further suggests that warming could undermine long-term predictability of water supply and complicate adaptation planning.
Despite these advances, several key uncertainties remain. Although we employed rigorous dataset screening, observational constraints remain due to sparse in-situ networks and satellite retrieval biases. The coarse spatial resolution of SM products may also underestimate localized heterogeneity, particularly in steep or high-relief regions [zeng2015evaluation, su2011tibetan]. Model simulations also exhibit known deficiencies in representing SM processes in high-altitude and semi-arid environments. Moreover, while Pp is a primary driver, the contribution of snow accumulation, glacier melt, and surface runoff to entropy dynamics remains insufficiently assessed. Finally, evidence of abrupt regime shifts is currently confined to a subset of models, leaving open the question of whether these signals represent genuine thresholds or model-specific artifacts.
@@
Figure 2: Evolution of $S_{\mathrm{TP}}$ and its relationship with ENSO. (a) Temporal evolution of $S_{\mathrm{TP}}$ from 2002–2024 (black), showing a significant decline with a breakpoint in 2018 identified by segmented regression. The decrease is weak before 2018 ($p<0.1$) but accelerates afterwards ($p<0.001$). (b) Mean SM before and after the 2018 breakpoint, illustrating that wetter conditions correspond to lower disorder. (c) $S_{\mathrm{TP}}$ normalized by grid number NN, showing that locations with higher mean SM ($\geq 0.15$) consistently exhibit lower entropy. (d) Joint evolution of $S_{\mathrm{TP}}$ (black) and $S_{\mathrm{ENSO}}$ (blue), with lag correlations peaking at 0.59 when $S_{\mathrm{TP}}$ lags $S_{\mathrm{ENSO}}$ by 6 months, highlighting ENSO modulation of SM entropy on the TP.
Figure 3: Relationships between the spatiotemporal patterns of soil moisture EM1, $S_{\mathrm{TP}}$ and ENSO. (a) Spatial patterns of the soil moisture EM1, highlighting a west–southwest dipole. (b) The 12-month smoothed temporal evolution of soil moisture EM1 (SM $V_{1}$, blue) and ENSO index (MEI.v2, red), showing a strong correlation (r=0.6r=-0.6, $p<0.001$). (c) Co-evolution of $S_{\mathrm{TP}}$ (blue) and 3-year running mean SM difference between negative and positive phase regions in SM $U_{1}$ ($\delta_{\mathrm{SM}}(U_{1}^{N}-U_{1}^{P})$, red), with a correlation of 0.66 ($p<0.001$), indicating that larger spatial contrasts correspond to higher entropy. (d) Violin plots of the two-month running mean SM anomaly difference between positive and negative EM1 phase regions, with individual data points overlaid, during El Niño (red), La Niña (blue), and neutral (gray) periods. Boxplots show the median, the 25th and 75th percentiles (box edges), 95% confidence interval (whiskers) and mean values (yellow triangles). Significant differences (***, $p<0.001$) were identified between El Niño and La Niña, and between La Niña and neutral, based on both one-way ANOVA and two-sided Mann–Whitney $U$-test, whereas no significant difference was found between El Niño and neutral.
Figure 4: Spatiotemporal patterns of precipitation EM1 and its relationships with SM and ENSO. (a) Spatial patterns of the precipitation EM1 (Pp $U_{1}$). (b) Temporal evolution of precipitation EM1 (Pp $V_{1}$, blue) and soil moisture EM1 (SM $V_{1}$, red), with a strong lagged correlation peaking when SM lags Pp by one month, indicating rapid SM response to precipitation variability in the TP. (c) Mean monthly Pp anomaly difference between ENSO phases (El Niño - La Niña). (d) Difference in vertically integrated water vapour flux anomaly (vectors) and its divergence anomaly (VIMDF, shaded) between ENSO phases (El Niño - La Niña). Asterisks denote regions with statistically significant differences based on a two-tailed Welch’s t-test ($p<0.1$).
@@ 4.1.2 Meteorological data
This study employs monthly meteorological variables including precipitation (Pp) and evapotranspiration (ET) from the GLDAS-NOAH dataset provided at $0.25^{\circ}\times 0.25^{\circ}$ resolution. Near-surface (1000-hPa) air temperature from ERA5 was used to assess the spatial covariability between ENSO signals and SM $V_{1}$. Monthly total column vertically integrated water vapour flux and its divergence from ERA5 were used to investigate ENSO-related moisture transport influencing Pp over the TP. ENSO variability was quantified using the Multivariate ENSO Index Version 2 (MEI.v2), which combines sea level pressure, sea surface temperature, low-level winds, and outgoing longwave radiation to provide a physically consistent representation of coupled ocean–atmosphere dynamics [zhang2019towards]. Since MEI.v2 does not define standardized thresholds for the onset or classification of ENSO events, the Oceanic Niño Index (ONI) was additionally used to identify and categorize ENSO phases. El Niño (La Niña) events are defined as periods when the ONI exceeds $0.5~^{\circ}\mathrm{C}$ (falls below $0.5~^{\circ}\mathrm{C}$) for at least five consecutive months(see Supplementary Fig. 20).
@@ 4.3 Eigen Microstates Entropy
which measures how evenly the system’s dynamics are distributed among the eigen microstates. Low $S_{\text{EM}}$ indicates that variability is concentrated in one or a few coherent modes, reflecting an ordered system with strong organization. High $S_{\text{EM}}$ reflects a more uniform distribution with many competing modes, corresponding to disordered or noisy states with no dominant spatial–temporal pattern [Liu2025Phase]. In the theoretical limit of a perfectly ordered system (e.g., condensation into a single dominant eigen microstates), $p_{1}\rightarrow 1$ and $S_{\text{EM}}\rightarrow 0$. Conversely, in a maximally disordered system (e.g., purely random system) with all eigen microstates contributing equally, $p_{I}\rightarrow 1/N$ and $S_{\text{EM}}\rightarrow\ln{N}$. Therefore, $S_{\text{EM}}$ serves as a sensitive and interpretable measure of disorder.

### v2

@@ Abstract
The Tibetan Plateau (TP), known as the “Asian Water Tower,” is currently undergoing a rapid wetting trend. While this moisture increase is often viewed as beneficial for water availability, it remains unclear whether the hydrological system itself is becoming more resilient or drifting toward instability. Here, we apply an entropy-based framework to quantify the changing structural organization of the TP’s soil moisture system. We show that from 2000 to 2024, regional wetting has driven a long-term decline in entropy, reflecting an increase in system order and stability due to enhanced hydrological buffering capacity. This stability is modulated by the El Niño–Southern Oscillation (ENSO), which regulates regional heterogeneity via a distinct spatial dipole. Crucially, however, CMIP6 climate projections reveal an alarming reversal: future warming triggers a rise in entropy. This transition signals a loss of systemic resilience, characterized by intensified spatial disorder and potential abrupt regime shifts by the mid-century. Our findings suggest that while current wetting provides a stabilizing buffer, continued warming is projected to amplify spatial heterogeneity, thereby destabilizing the Asian Water Tower, with significant risks for downstream water security.
@@ 1 Introduction
A further challenge lies in the lack of a theoretical framework capable of assessing the systemic behavior of SM beyond localized or pixel-based variability. SM dynamics arise from nonlinear interactions among Pp, ET, snowmelt and vegetation feedbacks, and are additionally influenced by stochastic climate variability [seneviratne2010investigating]. Under external perturbations such as sustained warming, these coupled processes may trigger amplifying mechanisms, such as plant-soil moisture positive feedback [hu2024plant], potentially pushing the system toward alternative regimes when stability decreases. Capturing such large-scale behavior requires a metric that goes beyond mean trends to quantify organization, heterogeneity and resilience. Entropy provides precisely such a diagnostic lens. By quantifying the degree of disorder, entropy reveals whether variability is concentrated in a few dominant modes (order) or dispersed across many (disorder), offering a physically interpretable measure of systemic structure and early-warning signals of abrupt transitions [vranken2015review, tirabassi2023entropy, meng2020complexity]. Here, we leverage recent advances in statistical physics to estimate entropy directly from observational and model data using an eigen microstate approach [hu2019condensation, liu2022renormalization, ma2024increased, zhang2024eigen, wang2024holistic, xie2025ecosystem]. This method extracts dominant system states to characterize their probability distribution [Liu2025Phase], offering a tractable framework for tracking SM disorder under both observed variability and projected climate warming.
@@ 2.1 Significant wetting over the TP during 2000–2024
Considering the inconsistencies across current SM datasets, we first identify the dataset most suitable for our analysis of the spatiotemporal dynamics of SM across the TP. We develop a robust evaluation framework that jointly considers temporal correlation, spectral characteristics, and autocorrelation structure, using in situ observations at 5 cm depth from the Tibet-Obs network, which comprises Maqu, Naqu, Ali, and Shiquanhe subregions (Supplementary Fig. 1). The evaluation incorporates three complementary metrics designed to assess different aspects of temporal fidelity. (1) temporal correlation ($R_{T}$; see Methods), which quantifies the degree of synchronicity between the evaluated products and site-level measurements. A high $R_{T}$ indicates strong agreement in the timing and fluctuation patterns of SM variability. (2) spectral correlation ($R_{F}$), calculated from the Fourier transforms of evaluated products and observational time series. $R_{F}$ assesses the preservation of frequency-dependent variability, ensuring the dataset reproduces not only first-order trends but also higher-order dynamical structures embedded in the observed record. (3) autocorrelation sequence similarity ($R_{ACF}$), based on lagged autocorrelation coefficients at lags 1–10 (AC1–AC10). $R_{ACF}$ captures the persistence and memory structure of SM dynamics, with higher similarity indicating a better capability of the evaluated product to replicate the temporal continuity and internal variability of real-world processes.
@@ 2.2 ENSO-Driven Modulation of SM Entropy on the TP
A broadly increasing trend in SM across the TP is reshaping regional hydrological dynamics by enhancing soil water storage, altering the partitioning of ET, and modifying runoff generation. These changes affect both the timing and magnitude of water availability across the Plateau and its downstream basins, with important implications for the stability of the regional water cycle. To evaluate soil water stability, we use eigen microstate entropy as a diagnostic of system disorder. This metric quantifies how SM variability is distributed across dominant spatial-temporal modes: lower entropy indicates a state that is governed by a few coherent modes, whereas higher entropy reflects a more disordered state lacking clear organization (see Methods). Applying this framework to the TP’s SM reveals a distinct long-term reorganization and clear teleconnections with ENSO events.
@@ 2.4 Projected drying and increasing disorder of the SM system over the TP
Rapid warming over the TP is expected to elicit nonlinear responses in the SM system, potentially triggering abrupt regional shifts under future climate scenarios [liu2023teleconnections, terpstra2025assessment, zaitchik2023wetting, qiao2023soil]. Projected increases in ENSO frequency and amplitude [cai2021changing] may further destabilize SM patterns through the teleconnections described above. Moreover, Soil moisture–atmosphere coupling over the TP may propagate via Rossby wave pathways, potentially modulating remote climate extremes, including heatwaves over Europe and East Asia [jiang2024dry, qiao2023soil], highlighting the importance of assessing projected TP soil moisture changes under warming scenarios. In complex systems, rising entropy has been proposed as an early-warning signal of reduced stability, preceding abrupt transitions across diverse contexts, from theoretical models to biological processes [tirabassi2023entropy, vranken2015review]. To illustrate this, we consider a Local Facilitation Cellular Automaton (LFCA) model [kefi2007local], which undergoes a bifurcation-like shift when a control parameter changes. Unlike the one-dimensional Ornstein-Uhlenbeck process often used in climate tipping point studies [smith2023reliability], the LFCA model is inherently high-dimensional, allowing it to capture interactions among multiple components. This property makes it more representative of real-world complexity and well-suited to eco-hydrological processes such as vegetation-wetness interactions [tirabassi2023entropy]. As shown in Fig. 5a, system state declines as the control parameter weakens, while the entropy of system state ($S_{\mathrm{LFCA}}$) rises in advance of the threshold, signaling the approach of an abrupt transition.
@@ 3 Discussion and Conclusion
Rising entropy and enhanced spatial heterogeneity across the TP suggest that water availability will become increasingly uneven. This may have cascading implications for major downstream river basins in South and East Asia, which collectively support nearly two billion people. Negative spring SM anomalies in the western TP, for instance, have been linked to extreme summer Pp in the Yangtze River Basin, intensifying downstream flood risk [zhu2023diagnosing]. At the same time, SM-driven circulation anomalies across the Plateau, reinforced by strong land–atmosphere feedbacks, are identified as key drivers of concurrent extremes, including East Asian heatwaves and South Asian floods [he2023extremely]. Increasing TP instability may therefore pose significant threats to agricultural production, hydropower generation, and water security across densely populated regions [yao2022imbalance, cui2023non]. The emergence of regime-shift behavior under high-emission scenarios further suggests that warming could undermine the long-term predictability of water supply and complicate adaptation planning.
Despite these advances, several key uncertainties remain. In this study, entropy is not interpreted as a direct measure of stability, but rather as a diagnostic of the system’s degree of structural organization. Rising entropy reflects a progressive dispersion of variability across multiple competing modes and is often interpreted as a loss of coherent organization associated with reduced resilience and heightened sensitivity to perturbations in complex, high-dimensional systems. [de2014measure, tirabassi2023entropy, Liu2025Phase]. Although we employed rigorous dataset screening, observational constraints remain due to sparse in-situ networks and satellite retrieval biases. The coarse spatial resolution of SM products may also underestimate localized heterogeneity, particularly in steep or high-relief regions [zeng2015evaluation, su2011tibetan]. Model simulations also exhibit known deficiencies in representing SM processes in high-altitude and semi-arid environments. Moreover, while Pp is a primary driver, the contribution of snow accumulation, glacier melt, and surface runoff to entropy dynamics remains insufficiently assessed. Finally, evidence of abrupt regime shifts is currently confined to a subset of models, leaving open the question of whether these signals represent genuine thresholds or model-specific artifacts.
@@
Figure 2: Evolution of $\boldsymbol{S_{\mathrm{TP}}}$ and its relationship with ENSO. (a) Temporal evolution of $S_{\mathrm{TP}}$ from 2002–2024 (black), showing a significant decline with a breakpoint in 2018 identified by segmented regression. The decrease is weak before 2018 ($p<0.1$) but accelerates afterwards ($p<0.001$). (b) Mean SM before and after the 2018 breakpoint, illustrating that wetter conditions correspond to lower disorder. (c) $S_{\mathrm{TP}}$ normalized by grid number NN, showing that locations with higher mean SM ($\geq 0.15$) consistently exhibit lower entropy. (d) Joint evolution of $S_{\mathrm{TP}}$ (black) and $S_{\mathrm{ENSO}}$ (blue), with lag correlations peaking at 0.59 when $S_{\mathrm{TP}}$ lags $S_{\mathrm{ENSO}}$ by 6 months, highlighting ENSO modulation of SM entropy on the TP.
Figure 3: Relationships between the spatiotemporal patterns of soil moisture EM1, $\boldsymbol{S_{\mathrm{TP}}}$ and ENSO. (a) Spatial patterns of the soil moisture EM1, highlighting a west–southwest dipole. (b) The 12-month smoothed temporal evolution of soil moisture EM1 (SM $V_{1}$, blue) and ENSO index (MEI.v2, red), showing a strong correlation (r=0.6r=-0.6, $p<0.001$). (c) Co-evolution of $S_{\mathrm{TP}}$ (blue) and 3-year running mean SM difference between negative and positive phase regions in SM $U_{1}$ ($\delta_{\mathrm{SM}}(U_{1}^{N}-U_{1}^{P})$, red), with a correlation of 0.66 ($p<0.001$), indicating that larger spatial contrasts correspond to higher entropy. (d) Violin plots of the two-month running mean SM anomaly difference between positive and negative EM1 phase regions, with individual data points overlaid, during El Niño (red), La Niña (blue), and neutral (gray) periods. Boxplots show the median, the 25th and 75th percentiles (box edges), 95% confidence interval (whiskers) and mean values (yellow triangles). Significant differences (***, $p<0.001$) were identified between El Niño and La Niña, and between La Niña and neutral, based on both one-way ANOVA and two-sided Mann–Whitney $U$-test, whereas no significant difference was found between El Niño and neutral.
Figure 4: Spatiotemporal patterns of precipitation EM1 and its relationships with SM and ENSO. (a) Spatial patterns of the precipitation EM1 (Pp $U_{1}$). (b) Temporal evolution of precipitation EM1 (Pp $V_{1}$, blue) and soil moisture EM1 (SM $V_{1}$, red), with a strong lagged correlation peaking when SM lags Pp by one month, indicating a rapid SM response to precipitation variability in the TP. (c) Mean monthly Pp anomaly difference between ENSO phases (El Niño - La Niña). (d) Difference in vertically integrated water vapour flux anomaly (vectors) and its divergence anomaly (VIMDF, shaded) between ENSO phases (El Niño - La Niña). Asterisks denote regions with statistically significant differences based on a two-tailed Welch’s t-test ($p<0.1$).
@@ 4.1.2 Meteorological data
This study employs monthly meteorological variables including precipitation (Pp) and evapotranspiration (ET) from the GLDAS-NOAH dataset provided at $0.25^{\circ}\times 0.25^{\circ}$ resolution. Near-surface (1000-hPa) air temperature from ERA5 was used to assess the spatial covariability between ENSO signals and SM $V_{1}$. The monthly total column vertically integrated water vapour flux and its divergence from ERA5 were used to investigate ENSO-related moisture transport influencing Pp over the TP. ENSO variability was quantified using the Multivariate ENSO Index Version 2 (MEI.v2), which combines sea level pressure, sea surface temperature, low-level winds, and outgoing longwave radiation to provide a physically consistent representation of coupled ocean–atmosphere dynamics [zhang2019towards]. Since MEI.v2 does not define standardized thresholds for the onset or classification of ENSO events, the Oceanic Niño Index (ONI) was additionally used to identify and categorize ENSO phases. El Niño (La Niña) events are defined as periods when the ONI exceeds $0.5~^{\circ}\mathrm{C}$ (falls below $0.5~^{\circ}\mathrm{C}$) for at least five consecutive months(see Supplementary Fig. 20).
@@ 4.3 Eigen Microstates Entropy
which measures how evenly the system’s dynamics are distributed among the eigen microstates. Low $S_{\text{EM}}$ indicates that variability is concentrated in one or a few coherent modes, reflecting an ordered system with strong organization. High $S_{\text{EM}}$ reflects a more uniform distribution with many competing modes, corresponding to disordered or noisy states with no dominant spatial–temporal pattern [Liu2025Phase]. In the theoretical limit of a perfectly ordered system (e.g., condensation into a single dominant eigen microstate), $p_{1}\rightarrow 1$ and $S_{\text{EM}}\rightarrow 0$. Conversely, in a maximally disordered system (e.g., purely random system) with all eigen microstates contributing equally, $p_{I}\rightarrow 1/N$ and $S_{\text{EM}}\rightarrow\ln{N}$. Therefore, $S_{\text{EM}}$ serves as a sensitive and interpretable measure of disorder.